Please help me with this.


5/6x + 1/3 = 3/4 x -1/2=??

Answer by stanbon(75887)

(5/6)x + 1/3 = (3/4)x -1/2
[(5/6)-(3/4)]x = -[1/2+1/3]
[2/24]x = -[5/6]
[1/12]x = -[5/6]
x = -[5/6]*12
x= -10
===========
Cheers,
Stan H.